# KANAZAWA-25-04 Phase structure analysis of 2d lattice CP(1) model with $\theta$ term using tensor renormalization group method

### v1

@@ 1 Introduction
The strong CP problem is an unresolved issue in particle physics [1, 2]. It arises from the difficulty of investigating the $\theta$ term in quantum chromo-dynamics (QCD). Since QCD exhibits non-perturbative phenomena such as confinement, many of its properties have been studied through numerical calculations based on lattice field theory combined with Monte Carlo (MC) methods. However, the presence of the $\theta$ term renders the associated Boltzmann weight a complex number, which leads to the sign problem in MC simulations and makes the investigation of this term particularly difficult [3].
@@
On the other hand, in the case of CP(1)($N=2$), there are qualitative predictions for the phase structure thanks to its relation to the Haldane conjecture. The conjecture states that the ground state of a one-dimensional spin-$s$ antiferromagnet drastically changes depending on the value of $s$, which is shown by using the effective model O(3) nonlinear $\sigma$ model [20, 21]. In this correspondence, the spin $s$ of the antiferromagnet maps to the $\theta$ parameter of the O(3) model, and it is found that the ground state is gapless at $\theta=\pi$, while it has a gap at $\theta=0$. When the ground state is gapless, the system is at a critical point where its universality class is reflected in the critical exponents. Analytical studies of the O(3) model using non-Abelian bosonization, etc., [22, 23] have shown that its universality class is SU(2)k=1 Wess-Zumino-Witten (WZW) model [24, 25]. Numerical studies of the lattice O(3) model have further confirmed the existence of the corresponding critical point at $\theta=\pi$ [26, 27, 28]. Since the CP(1) model and the O(3) nonlinear $\sigma$ model are equivalent in the continuum limit, it is expected that the lattice CP(1) model also exhibits the same criticality at $\theta=\pi$.
@@
The rest of the paper is organized as follows. In section 2, we fix the notations used throughout the paper. The action and the partition function of CP(1) model on the square lattice are defined in this section. In section 3, we present a tensor network representation of the partition function of the model, and examine the accuracy of the initial tensor. Numerical results for the phase structure analysis, including CFT data, are shown in section 4. Our concluding remarks are given in section 5. Appendix A provides the details of constructing the initial tensor using the character expansion. In appendix B, we describe the quadrature method on hyperspheres. Appendix C presents the exact result of the free energy on a $2\times 2$ lattice.
@@ 4.2 Degeneracy of ground state
Figure 6: Grand-state degeneracy $X$ in Eq. (27) as a function of $\beta$ at $\theta=\pi$ on $V=4^{2}$$4^{7}$.
@@
Figure 6 shows $X$ as a function of $\beta$ at $\theta=\pi$. For $\beta<0.5$, we find that $X=2$ persists even for sufficiently large volumes, thereby, confirming the presence of the first-order phase transition in the low $\beta$ region. In contrast, for $\beta>0.5$, $X$ takes non-integer values, and thus the presence or absence of ground-state degeneracy cannot be unambiguously determined. It also remains unclear at which point the first-order phase transition terminates. On the other hand, Fig. 7 displays $X$ for the case $\theta\neq\pi$, where we consistently obtain $X=1$ for all $\beta$ in sufficiently large volumes. This result clearly indicates that no first-order phase transition occurs away from $\theta=\pi$ within the $\beta$ region we investigated.
@@ 4.3 Analysis based on CFT data
In the present work, we also adopt the level spectroscopy method to investigate the phase structure of the lattice CP(1) model, which is equivalent to $O(3)$ nonlinear $\sigma$ model in the continuum limit. However, note that while the $O(3)$ model possesses a global SU(2) symmetry, the XY model does not, and consequently the perturbations allowed around the corresponding CFT differ between the two cases. Therefore, for a while, let us briefly discuss the perturbation for the case of the $O(3)$ model. The form of the perturbation is predicted by combining the relation between antiferromagnets and $O(3)$ known as the Haldane conjecture with the application of bosonization to antiferromagnets. Figure 10 schematically represents this prediction. The arrow (a) in Fig. 10 means that the $O(3)$ nonlinear $\sigma$ model at $\theta=\pi$ is the low-energy effective field theory of a spin-$\frac{1}{2}$ Heisenberg model with an additional next-nearest-neighbor interaction, in the regime where the next-nearest-neighbor coupling is sufficiently large [42, 22]. In this correspondence, the coupling constant $J_{2}$ of the next-nearest-neighbor interaction in the Heisenberg model maps onto the inverse coupling $\beta$ of the $O(3)$ model. On the other hand, in such large $J_{2}$ case, the Heisenberg model corresponds to the SU(2)k=1 WZW model with an additional marginal operator in the low-energy limit as shown in the arrow (b) in Fig. 10. This correspondence is revealed by non-Abelian bosonization analyses [43, 22], and note that $J_{2}$ maps onto $\lambda$, the coupling constant associated with the marginal operator. From the above arguments (a) and (b), the relation indicated by the arrow (c) in Fig. 10 can be predicted, i.e.,
@@
where the marginal coupling constant $\lambda$ are related with $\beta$ in the O(3) model as $\lambda=a(\beta-\beta_{\rm c})$ with a proportional constant $a$. In the equation above, the symbol $\sim$ indicates that irrelevant contributions are ignored (these contributions will be discussed later). Here, $S_{\mathrm{SU(2)}_{k=1}}$ has the chiral symmetry $\mathrm{SU(2)}_{\rm L}\times\mathrm{SU(2)}_{\rm R}$, and it is known that the lowest scaling dimension is $\frac{1}{2}$, which comes with a fourfold degeneracy. On the other hand, $S_{\rm marginal}$ preserves only the diagonal subgroup of the chiral symmetry, $\mathrm{SU(2)}$, and consequently the fourfold degeneracy is lifted into a singlet and a triplet once $\beta$ is shifted away from $\beta_{\rm c}$. This is a crucial distinction between the XY model and the $O(3)$ model. In the end, perturbation calculations for the SU(2)k=1 WZW model [44, 45] yield the scaling dimensions of the singlet $x_{\rm s}$ and triplet $x_{\rm t}$ as

### v2

@@ 1 Introduction
Quantum Chromodynamics (QCD) allows the presence of a topological $\theta$ term in its Lagrangian [1]. The presence of the $\theta$ term, however, renders the associated Boltzmann weight a complex number, which leads to the sign problem in MC simulations and makes the investigation of this term particularly difficult [2]. In nature, the value of the $\theta$ parameter is known to be extremely small, which is referred to as the strong CP problem [3, 4]. Despite this, investigating QCD with a finite $\theta$ term remains of theoretical interest, as it can help us better understand nonperturbative and topological aspects of the strong interaction.
@@
On the other hand, in the case of CP(1)($N=2$), there are qualitative predictions for the phase structure thanks to its relation to the Haldane conjecture. The conjecture states that the ground state of a one-dimensional spin-$s$ antiferromagnet drastically changes depending on the value of $s$, which is shown by using the effective model O(3) nonlinear $\sigma$ model [21, 22]. In this correspondence, the spin $s$ of the antiferromagnet maps to the $\theta$ parameter of the O(3) model, and it is found that the ground state is gapless at $\theta=\pi$, while it has a gap at $\theta=0$. When the ground state is gapless, the system is at a critical point where its universality class is reflected in the critical exponents. Analytical studies of the O(3) model using non-Abelian bosonization, etc., [23, 24] have shown that its universality class is SU(2)k=1 Wess-Zumino-Witten (WZW) model [25, 26]. Numerical studies of the lattice O(3) model have further confirmed the existence of the corresponding critical point at $\theta=\pi$ [27, 28, 29]. Furthermore, the study using D-theory regularization also produced results consistent with this prediction [30]. Since the CP(1) model and the O(3) nonlinear $\sigma$ model are equivalent in the continuum limit, it is expected that the lattice CP(1) model also exhibits the same criticality at $\theta=\pi$. In fact, numerical calculation, which employs a sign-problem-free formulation of the CP(1) model based on the Villain-type regularization of the U(1) gauge theory coupled to 2-flavor Higgs fields, has found the expected critical points [31].
@@
The rest of the paper is organized as follows. In section 2, we fix the notations used throughout the paper. The action and the partition function of CP(1) model on the square lattice are defined in this section. In section 3, we present a tensor network representation of the partition function of the model, and examine the accuracy of the initial tensor. Numerical results for the phase structure analysis, including CFT data, are shown in section 4. Our concluding remarks are given in section 5. Appendix A provides the details of constructing the initial tensor using the character expansion. In appendix B, we describe the quadrature method on hyperspheres. Appendix C presents the exact result of the free energy on a $2\times 2$ lattice. In appendix D, we discuss to what extent the important symmetries are maintained in numerical calculations.
@@ 4.2 Degeneracy of ground state
Figure 6: Ground-state degeneracy $X$ in Eq. (27) as a function of $\beta$ at $\theta=\pi$ on $V=4^{2}$$4^{7}$. The pink line represents the value for SU(2)k=1 WZW model in Eq. (31).
@@
Figure 6 shows $X$ as a function of $\beta$ at $\theta=\pi$. For $\beta<0.5$, we find that $X=2$ persists even for sufficiently large volumes, thereby, confirming the presence of the first-order phase transition in the low $\beta$ region. In contrast, for $\beta>0.5$, $X$ takes non-integer values, and thus the presence or absence of ground-state degeneracy cannot be unambiguously determined. When the system is tuned to its critical point, the value of $X$ can be represented using the partition function of the corresponding CFT. The modular invariant partition function for SU(2)k=1 WZW model on torus [41, 42] is known to be
$Z(\tau)=\left|\frac{1}{\eta}\sum_{n\in\mathbb{Z}}q^{n^{2}}\right|^{2}+\left|\frac{1}{\eta}\sum_{n\in\mathbb{Z}}q^{(n+\frac{1}{2})^{2}}\right|^{2},\hskip 14.22636ptq=e^{2\pi i\tau}$ (29)
where $\tau\in\mathbb{C}$ is the modular parameter and $\eta$ is the Dedekind $\eta$ function defined by
$\eta(\tau)=q^{\frac{1}{24}}\prod_{n=1}^{\infty}(1-q^{n}).$ (30)
In terms of the partition function, the value of $X$ for the SU(2)k=1 WZW model is given by
$X=\frac{Z(\tau=i)^{2}}{Z(\tau=2i)}=1.39324280(1).$ (31)
This value is shown as the pink line in Fig. 6, and the numerical data for $X$ approaches this line as the volume is increased for $\beta=0.55\text{ -- }0.6$. This behavior suggests that the critical point associated with the SU(2)k=1 WZW model lies within this range of $\beta$. To determine the critical point more accurately, we will perform a more detailed analysis using CFT information in the next subsection.
On the other hand, Fig. 7 displays $X$ for the case $\theta\neq\pi$, where we consistently obtain $X=1$ for all $\beta$ in sufficiently large volumes. This result clearly indicates that no first-order phase transition occurs away from $\theta=\pi$ within the $\beta$ region we investigated.
@@ 4.3 Analysis based on CFT data
In the present work, we also adopt the level spectroscopy method to investigate the phase structure of the lattice CP(1) model, which is equivalent to O(3) nonlinear $\sigma$ model in the continuum limit. However, note that while the O(3) model possesses a global SO(3) symmetry, the XY model does not, and consequently the perturbations allowed around the corresponding CFT differ between the two cases. Therefore, for a while, let us briefly discuss the perturbation for the case of the O(3) model. The form of the perturbation is predicted by combining the relation between antiferromagnets and O(3) known as the Haldane conjecture with the application of bosonization to antiferromagnets. Figure 10 schematically represents this prediction. The arrow (a) in Fig. 10 means that the O(3) nonlinear $\sigma$ model at $\theta=\pi$ is the low-energy effective field theory of a spin-$\frac{1}{2}$ Heisenberg model with an additional next-nearest-neighbor interaction, in the regime where the next-nearest-neighbor coupling is sufficiently large [47, 23]. In this correspondence, the coupling constant $J_{2}$ of the next-nearest-neighbor interaction in the Heisenberg model maps onto the inverse coupling $\beta$ of the O(3) model. On the other hand, in such large $J_{2}$ case, the Heisenberg model corresponds to the SU(2)k=1 WZW model with an additional marginal operator in the low-energy limit as shown in the arrow (b) in Fig. 10. This correspondence is revealed by non-Abelian bosonization analyses [48, 23], and note that $J_{2}$ maps onto $\lambda$, the coupling constant associated with the marginal operator. From the above arguments (a) and (b), the relation indicated by the arrow (c) in Fig. 10 can be predicted, i.e.,
@@
where the marginal coupling constant $\lambda$ are related with $\beta$ in the O(3) model as $\lambda=a(\beta-\beta_{\rm c})$ with a proportional constant $a$. In the equation above, the symbol $\sim$ indicates that irrelevant contributions are ignored (these contributions will be discussed later). Here, $S_{\mathrm{SU(2)}_{k=1}}$ has the chiral symmetry $\mathrm{SU(2)}_{\rm L}\times\mathrm{SU(2)}_{\rm R}$, and it is known that the lowest scaling dimension is $\frac{1}{2}$, which comes with a fourfold degeneracy. On the other hand, $S_{\rm marginal}$ preserves only the diagonal subgroup of the chiral symmetry, [^4] In Appendix D, we investigate to what extent the SO(3) symmetry and the CP symmetry at $\theta=\pi$ are maintained in our numerical calculation. SO(3), and consequently the fourfold degeneracy is lifted into a singlet and a triplet once $\beta$ is shifted away from $\beta_{\rm c}$. This is a crucial distinction between the XY model and the O(3) model. In the end, perturbation calculations for the SU(2)k=1 WZW model [49, 50] yield the scaling dimensions of the singlet $x_{\rm s}$ and triplet $x_{\rm t}$ as
@@ Appendix C Exact results for $2\times 2$ lattice
Appendix D Numerical investigation of SO(3) and CP symmetry breaking
We assess the extent to which the SO(3) symmetry and the CP symmetry at $\theta=\pi$ are violated due to the quadrature method and the coarse-graining procedure. To this end we compute the corresponding order parameters for each symmetry. In the following numerical results, the parameters for the initial tensor, and the coarse-graining algorithm are identical to those used in the main text.
The order parameter associated with the SO(3) symmetry may be defined as
$x_{i}:=\braket{z^{\dagger}\sigma_{i}z},$ (108)
where $z$ is the complex scalar field and $\sigma_{i}$ ($i=1,2,3$) denote the Pauli matrices. To evaluate this observable, we employ the impurity tensor network technique[55]. Figure 20 shows the numerical results for $x_{1}=\braket{z^{\dagger}\sigma_{1}z}$. For all values of $\beta$, $\theta$, and the volume $V$ shown in the figure, the SO(3) breaking remains below $O(10^{-12})$ and exhibits negligible volume dependence. We therefore conclude that the violation of the SO(3) symmetry is well controlled, and no strongly relevant perturbation is induced by the numerical procedures.
Figure 20: The order parameter for the SO(3) symmetry, $x_{1}$, as a function of $\beta$ for $V=4^{2}-4^{5}$.
As an order parameter for CP symmetry, we consider the topological charge density
$q:=\frac{\partial f}{\partial\theta},$ (109)
where $f$ is the free energy given in Eq. (26). We evaluate $q$ using the impurity method as well. Figure 21 presents the numerical results for $q$ at $\theta=\pi$ in the region $\beta_{\rm c}\lesssim\beta$, where CP symmetry implies that $q$ should vanish. Our results show that $q$ remains smaller than $O(10^{-5})$. Although this breaking appears larger than that observed for the SO(3) symmetry, the volume dependence around $\beta_{\rm c}\approx 0.6$ is weak. Thus the CP symmetry is expected to be under sufficient control within our numerical setup.
Figure 21: The order parameter for the CP symmetry, the topological charge density $q$, as a function of $\beta$ for $V=4^{2}-4^{5}$ at $\theta=\pi$.